OpenStudy (anonymous):

Tonight is the end of the semester and I need a little help! I'm not sure how to do anything with Linear Equations and Inequalities. I've tried to learn this many times but I don't understand any of it. Can someone please help me out?

OpenStudy (anonymous):

study your textbook always helps. Works fine for me

OpenStudy (anonymous):

I've tried I just don't understand it.

OpenStudy (anonymous):

try to internet. Youtube is good

OpenStudy (anonymous):

I have. I just need a one-on-one with someone teaching me.

OpenStudy (phi):

you could post a question that you would like to know how to solve

OpenStudy (anonymous):

I don't have any particular questions. I basically just need to know how to solve and/or graph them.

OpenStudy (phi):

example ?

OpenStudy (anonymous):

Yes!

OpenStudy (anonymous):

And slopes. For example: Find the slope of the line assing through the points (-3, 7) and (1, 2).

OpenStudy (anonymous):

passing*

OpenStudy (phi):

search Khan's site for anything you might want to review. He has lots on slope. For example http://www.khanacademy.org/math/algebra/linear-equations-and-inequalitie/slope-and-intercepts/v/slope-of-a-line-2

OpenStudy (anonymous):

A linear equation is one with the variables having no power other than 1, such as x and y. If something, y, grows proportional to x, such as y=mx, that is a simple linear equation. If y is not 0 but b when x = 0, we get the general form of a straight line: y=mx + b. The repairman charges $50 per hour (h) and $25 for showing up at all: Cost = ($50/hr)(h) + $25. A two-hour job (h=2) will cost you $125. If you have two such lines y1 = m1 x + b1 y2 = m2 x + b2, they will intersect somewhere unless parallel (m1=m2, same slopes) To "solve" that pair of equations is to find where they meet x = (b2 - b1)/(m1-m2) where you know all these values but x. y is obtained by substituting this x into one of your equations, y1 or y2. There are lots of other interesting aspects, but these are often most useful.